# Green Finance and Carbon Emissions: A Nonlinear and Interaction Analysis Using Bayesian Additive Regression Trees

### v1

@@ 5.3.2 Interaction effects between GFI and TEC
Figure 6 presents SHAP interaction effect plots between TEC and GFI, aiming to gain insights into complex feature interactions and uncover latent patterns. The left panel shows that when TEC is relatively low (below 10000) and GFI is also at a low level, the SHAP values are predominantly positive. This indicates that under conditions of low energy consumption and limited green finance development, carbon emission intensity tends to increase. As TEC increases beyond this level, accompanied by rising GFI levels, the SHAP values exhibit a clear downward trend and eventually turn negative. This suggests that when energy consumption is high, the interaction with higher levels of green finance leads to a negative marginal effect on CEI. In other words, the expansion of green finance may help decouple energy consumption from emissions, alleviating the environmental pressure typically associated with high energy use. In the right panel, as both TEC and GFI increase, the SHAP values shift from yellow to purple. This color gradient visually reinforces the finding that the marginal harm of energy consumption on carbon intensity becomes weaker, or even beneficial under high GFI conditions.
Taken together, the interaction between GFI and TEC reveals a clear nonlinear and synergistic relationship in shaping carbon emission intensity. Green finance not only directly influences CEI but also plays a critical moderating role by mitigating the environmental cost of energy use. These findings emphasize the importance of integrating financial instruments into carbon mitigation strategies. In particular, energy-intensive regions stand to benefit more from targeted green finance policies that can curb emissions even amid rising energy demand. Moreover, accounting for such interaction effects contributes to a more nuanced understanding of how various policy instruments, including financial incentives, energy regulation and regional planning, can be effectively coordinated to achieve long-term environmental objectives.
In summary, the interaction between population and GFI in shaping carbon emission intensity is highly complex and stage-dependent. These findings highlight the need for regionally differentiated green finance policies and call for particular attention to the potential “saturation effect" of green finance in megacities.
@@
Figure 6: SHAP interaction plots between GFI and TEC. The left panel shows the SHAP values of TEC across varying levels of GFI. The x-axis represents TEC, and each point is colored according to the GFI level, with lighter colors (yellow) indicating higher levels of green finance development. The y-axis represents the SHAP value of TEC, reflecting its marginal contribution to predicted CEI. The right panel provides a complementary visualization, where the x-axis shows TEC and the y-axis shows GFI. Each point represents an observation, with the color indicating the sum of their SHAP values on CEI (the contribution jointly made by the two variables). Yellow tones denote a stronger positive contribution to CEI, while purple tones indicate a negative or suppressing effect. This visualization highlights the combined marginal effects of the two features rather than their pure interaction term.
@@ 6 Conclusion
This study employs BART to investigate the complex relationships between green finance and carbon emissions across 30 Chinese provinces from 2000 to 2023. Through careful model tuning, evaluation and diagnostic testing, we confirm the strong predictive power and robustness of the BART approach in modeling environmental-economic interactions. The tuned model achieves high accuracy and generalizability, making it a reliable basis for further interpretation. Our analysis yields several key findings: (1) green finance is a relevant predictor of carbon emissions, but with a non-linear pattern. The impact of GFI on CEI exhibits a distinct inverted U-shaped pattern. In early stages of green finance development, CEI may increase due to initial resource inputs. However, once GFI surpasses a certain threshold around 0.3, it begins to significantly reduce carbon intensity. This pattern is consist with the findings of Wu et al. wu2024impact , and underscores the importance of long-term commitment and maturity in green financial systems; (2) regional heterogeneity is significant. The effects of GFI on CEI vary by region. Western regions benefit from early green finance adoption even at low levels, while Eastern provinces, where GFI is more mature. This indicates the strongest emission-reducing effects in Eastern provinces. Central China appears to be in a transition phase. These differences point to the necessity of region-specific green finance policies; (3) Green finance interacts with energy consumption, jointly influencing the outcome. Interaction analysis reveals that GFI moderates the effects of energy consumption on CEI. Specifically, high GFI can offset the emissions typically associated with high TEC, reflecting that green finance, under certain conditions, can weaken the positive correlation between energy consumption and carbon emissions, thereby playing a moderating role in emission reduction; (4) CPRI is not statistically significant. Contrary to expectations, CPRI does not exhibit a statistically significant impact on CEI in our model. However, this finding should not be interpreted as evidence that climate risk is irrelevant to carbon emissions. Rather it may stem from the limitations of using annual, province-level data, which may not adequately capture the more nuanced or delayed effects of climate risk. Climate risk often exerts its influence with a temporal lag and is likely to have a more pronounced impact at the micro level, particularly on individual firms or sectors. This highlights the need for future research that utilizes more granular data and incorporates potential lag structures to better assess the complex relationship between climate risk and emissions.
@@
These findings suggest that while green finance is a powerful tool for reducing carbon emissions, its design and implementation must account for nonlinear effects, interaction dynamics and regional heterogeneity. In particular, in highly urbanized areas, attention should be paid to the diminishing returns of green finance, which may necessitate complementary policies, such as technological innovation or integrated urban planning. Based on these findings, the analysis can be further extended to incorporate a spatial dimension by using geographical information across provinces to examine whether spatial spillover effects or regional clustering patterns influence the effectiveness of green finance policies and carbon reduction outcomes.

### v2

@@ 5.4.1 Interaction effects between GFI and TEC
Figure 6 presents the SHAP interaction effects between total energy consumption (TEC) and green finance (GFI), illustrating how their joint variations influence carbon emission intensity (CEI). It can be observed that when both TEC and GFI are relatively low, the SHAP values are predominantly positive, indicating that limited energy use and underdeveloped green finance are associated with higher CEI levels. As TEC and GFI increase, the SHAP values exhibit a downward trend, eventually even turning negative. It should be noted, however, that this decline in SHAP values does not imply that higher energy consumption directly reduces carbon intensity.
This phenomenon can be explained by the definition of CEI and the potential role of green finance. In this study, CEI is defined as carbon emissions per unit of GDP, and higher TEC is generally accompanied by increased economic output, which expands the GDP denominator and leads to a lower measured CEI. High levels of green finance may promote improvements in energy efficiency and transitions to cleaner energy sources, thereby weakening the marginal impact of energy consumption on CEI. Moreover, in more advanced stages of green finance development, economic scale expansion may offset part of the emission-promoting effect of increased energy consumption, further reducing carbon intensity per unit of output. In other words, the observed decline in SHAP values reflects the combined moderating effects of green finance and economic growth on carbon intensity, rather than a direct emission-reducing effect of energy consumption.
@@
Figure 6: SHAP interaction plots between GFI and TEC. The plot shows the SHAP values of TEC across varying levels of GFI. The x-axis represents TEC, and each point is colored according to the GFI level, with lighter colors (yellow) indicating higher levels of green finance development. The y-axis represents the SHAP value of TEC, reflecting its marginal contribution to predicted CEI. This visualization highlights the combined marginal effects of the two features rather than their pure interaction term.
Overall, the interaction between GFI and TEC reveals a clear nonlinear and synergistic relationship in shaping carbon emission intensity. Specifically, green finance not only directly influences CEI but also plays a critical moderating role by weakening the marginal effect of energy consumption on carbon intensity. These findings emphasize the importance of integrating financial instruments into carbon mitigation strategies. In particular, energy-intensive regions stand to benefit more from targeted green finance policies that can curb emissions even amid rising energy demand. Moreover, accounting for such interaction effects, including TEC × GFI, contributes to a more nuanced understanding of how various policy instruments, such as financial incentives, energy regulation, and regional planning, can be effectively coordinated to achieve long-term environmental objectives.
@@ 6 Conclusion
This study employs BART to investigate the complex relationships between green finance and carbon emissions across 30 Chinese provinces from 2000 to 2022. Through careful model tuning, evaluation and diagnostic testing, we confirm the strong predictive power and robustness of the BART approach in modeling environmental-economic interactions. The tuned model achieves high accuracy and generalizability, making it a reliable basis for further interpretation. Our analysis yields several key findings: (1) green finance is a relevant predictor of carbon emissions, but with a non-linear pattern. The impact of GFI on CEI exhibits a distinct inverted U-shaped pattern. This result generally aligns with the previous findings of wu2024impact . At low levels of GFI, the observed positive relationship with CEI likely reflects confounding structural factors rather than a direct effect of green finance. Once GFI surpasses a threshold of approximately 0.3, its genuine emission-reduction effect emerges, highlighting the importance of advanced development and long-term commitment in green financial systems. (2) Regional heterogeneity is apparent. The effects of GFI on CEI vary across regions. Eastern provinces, where green finance has a stronger presence, show the most pronounced carbon reduction effects at higher GFI levels. Central and Western regions, in contrast, remain in a transitional stage, with the mitigation potential of green finance not yet fully realized. These differences emphasize the need for region-specific strategies that account for local economic structures, technological capabilities, and policy maturity. (3) Green finance interacts with energy consumption, jointly influencing the outcome. GFI moderates the effects of TEC on CEI. High GFI can offset the positive emissions effects typically associated with high TEC, reflecting that green finance can weaken the correlation between energy consumption and carbon emissions under certain conditions. This underscores the importance of integrating financial instruments into carbon mitigation strategies, particularly in energy-intensive regions; (4) CPRI is not statistically significant. Contrary to expectations, CPRI does not exhibit a statistically significant impact on CEI in our model. However, this finding should not be interpreted as evidence that climate risk is irrelevant to carbon emissions. Rather, it may stem from the limitations of using annual, province-level data, which may not adequately capture the more nuanced or delayed effects of climate risk. Climate risk often exerts its influence with a temporal lag and is likely to have a more pronounced impact at the micro level, particularly on individual firms or sectors. This highlights the need for future research that utilizes more granular data and incorporates potential lag structures to better assess the complex relationship between climate risk and emissions.
@@
These findings suggest that while green finance is a powerful tool for reducing carbon emissions, its design and implementation must account for nonlinear effects, interaction dynamics and regional heterogeneity. Based on these findings, the analysis can be further extended to incorporate a spatial dimension by using geographical information across provinces to examine whether spatial spillover effects or regional clustering patterns influence the effectiveness of green finance policies and carbon reduction outcomes.